If 6 identical articles can be bought for rupees 2 whole number 6 upon 17 . Find the cost of each article...

Grade:7

2 Answers

swarup kumar majee
6 years ago
6 identical artical costs Rs 2 6/17=40/17
the cost of one article should be 40/17 divided by 6
so the cost of each article is Rs 20/17
Nitesh
6 years ago
Since 6 article costs 40/17 rupee So 1 article costs 40/17 ×1/6=20/51 rupeeSo the cost of each article is 20/51 rupees.